ROTATIONALLY-RESOLVED INFRARED SPECTROSCOPY OF CH_{3}F ISOLATED IN SOLID PARAHY-DROGEN

\underline{KAZUKI YOSHIOKA}, SHAWN PALMER, and DAVID T. ANDERSON, Depertment of Chemistry, University of Wyoming, Laramie, WY 82071-3838.

The high-resolution infrared spectrum of CH_{3}F isolated in solid parahydrogen is studied via FT-IR methods. The CH_{3}F doped parahydrogen crystals are prepared using the rapid vapor deposition method developed by Fajardo and co-workers. CH_{3}F was studied as a prototypical symmetric top molecule in a quantum cavity. The infrared spectroscopy suggests CH_{3}F undergoes nearly free rotation indicating it occupies a single substitution site. Spectroscopic results and analysis on the CH_{3}F rovibrational dynamics and nuclear spin relaxation will be presented.